# On the Compositional Generalization of Multimodal LLMs for Medical Imaging

### v1

@@ A.2 Other MLLMs (RQ 5)
The experimental results presented in Table 7 and Table 8 demonstrate that the CG persists across these different backbone architectures. This observation indicates that CG is not confined to a specific model type, thereby highlighting its universal applicability and robustness across various model frameworks. Such findings underscore the versatility of CG, suggesting it can be effectively integrated into a wide range of models.
@@ B.3 Data composition and Open-source Specification
| 4 | COVID CT COVID-CT | COVID19, Classification Dataset. | Zhao et al. (2020) |
@@
| 11 | PED-UFES-20 | Skin Multi Classification | Pacheco et al. (2020) |
@@
| 29 | APTOS 2019 Blindness | Blindness Level Identification 0 4 | Karthik et al. (2019) |
| 30 | DRIMBD | Quality Testing of Retinal Images | Prentasic et al. (2013) |
@@
| 35 | Blood Cell Images | Blood Cell Classificaion (Multi) | Mooney (2017) |
@@
| 37 | Chaoyang | Multi Classification of pathologists | Zhu et al. |
@@
| 54 | Shenzhen Chest X-Ray Set | COVID19, Classification Dataset. | Jaeger et al. (2014) |
@@
| 70 | HSA-NRL | Multi Classification of pathologists | Zhu et al. (2021) |
@@
| 84 | Human Protein Atlas | Multi-label Classification (Only green) | Le et al. (2022) |

### v2

@@ A.4 CG with All MAT-Triplet Elements from Different Sources
| Related Combination | Target Subset | HuatuoGPT | ||||
A.6 Details of Section 6: Exploring CG across Detection and Classification
The two backbones were trained separately in this experiment. For Next-Chat, we directly trained the model in its second training stage and fine-tuned it for 2 epochs with a learning rate of 2e-5, keeping all other training parameters at their default settings. Similarly, for MiniGPT-v2, we trained the backbone model from the second stage, starting with a learning rate of 2e-5 and gradually reducing it to 2e-6 over 3 epochs.
A.7 CG with Medical Multimodal LLM
In previous experiments, general MLLMs are selected to prevent the MLLM’s inherent medical knowledge from affecting CG results. Our experiments focus on how MLLMs leverage CG to interpret unseen medical images. If the model has learned some fundamental elements of the Target data, it would compromise the fairness of the experiments.
To demonstrate that our results still work on medical LLMs, we employed the same data combinations from Section 3.3 to investigate CG on medical MLLMs (we selected HuatuoGPT-Vision as the baseline).
The results in Table 15 demonstrate that the medical-expert MLLM can still leverage CG to enhance their performance on novel tasks, further supporting the validity and consistency of our findings.
@@ B.3 Data composition and Open-source Specification
B.4 Data Sources and Distribution
All Med-MAT data are sourced from public medical image challenges or widely used, high-impact datasets previously applied in deep learning training, ensuring reliable annotations. Before inclusion in Med-MAT, all datasets underwent label averaging where possible; test sets, in particular, were strictly balanced to ensure accuracy reliably reflects model performance. Each Med-MAT training subset contains 3,000 samples, while test sets maximize size under label balance constraints.
Appendix C Bad cases analysis and solutions
C.1 Bad case analysis
Some Trained models show minimal gains or even performance declines in Table 3, with classification accuracy lower than either the Baseline or Baseline+. After a thorough examination, we found that these Target datasets require more fine-grained medical condition classification. Beyond disease presence, they need detailed assessments, such as severity grading (e.g., bone age estimation, cancer staging) or distinguishing similar conditions (e.g., differentiating COVID-19 from pneumonia).
The Related combinations lack suitable fundamental elements: For CG, the training data must include the Target task’s core elements. Here, we use other "level classification/grading" tasks for generalization, but their criteria differ significantly, misaligning with the Target task’s needs.
Without defined grading standards, MLLMs lacking relevant knowledge can’t perform fine-grained tasks: Tasks like bone age assessment and cancer staging vary by criteria, and without this knowledge, MLLMs can’t accurately classify them.
C.2 Possible solutions
Few-shot prompting
As we illustrated before, most of the bad cases involve fine-grained tasks needing specialized knowledge. So, in order to minimize the effect of a lack of relevant knowledge, we also conducted few-shot experiments to add some target images in the prompts. Subset X-ray, Lung, Normal-COVID-Pneumonia was chosen for its simple structure, with LLaVA as the baseline. We randomly sampled n images per label for n-shot inference and repeated each experiment 3 times.
| Model | 0-shot | 2-shot | 3-shot | 4-shot |
| LLaVA | ||||
| LLaVA + CG |
The results in Table 16 demonstrate that training with CG combinations can improve the few-shot performance of MLLMs on downstream tasks, even when direct CG generalization is not effective.
Adding some Target data in training
As described in Section 5.2, we selected cases where CG alone couldn’t achieve satisfactory results and augmented their training sets with target data. The results in this section indicate that while CG may not directly enhance generalization, it accelerates the model’s adaptation to downstream tasks.
@@ C.2 Possible solutions
| 4 | COVID CT COVID-CT | COVID19, Classification Dataset | Zhao et al. (2020) |
@@
| 11 | PAD-UFES-20 | Skin Multi Classification | Pacheco et al. (2020) |
@@
| 29 | APTOS 2019 Blindness | Blindness Level Identification | Karthik et al. (2019) |
| 30 | DRIMDB | Quality Testing of Retinal Images | Prentasic et al. (2013) |
@@
| 35 | Blood Cell Images | Blood Cell Classificaion | Mooney (2017) |
@@
| 37 | Chaoyang | Multi Classification of pathologists | Zhu et al. (2021a) |
@@
| 54 | Shenzhen Chest X-Ray Set | COVID19, Classification Dataset | Jaeger et al. (2014) |
@@
| 70 | HSA-NRL | Multi Classification of pathologists | Zhu et al. (2021b) |
@@
| 84 | Human Protein Atlas | Multi-label Classification | Le et al. (2022) |